OpenStudy (erich):

(x-2)/3 - (x+2)/4 = 7/12 Solve for X

OpenStudy (anonymous):

(x-2)/3 - (x+2)/4 = 7/12 ((x-2)*4-(x+2)*3)/12=7/12 4x-8-3x-6=7 x=21

OpenStudy (anonymous):

1st, put all on same denominator: 4(x-2)/12-3(x+2)/12=7/12 Multiply all by 12: 4x-8-3x-6=7 <=> x-14=7 <=>x=21